# Transferring disentangled representations: bridging the gap between synthetic and real images

### v1

@@ 4 Conclusions
In this paper, we discussed the potential of transferring a Disentangled Representation learnt from a Source Dataset in a weakly supervised manner to a Target Dataset as a strategy to address disentanglement in real data, where supervision on the FoVs may be difficult or impossible. We identified three main scientific questions, summarised in Section 3.2, which we use to draw conclusions on our study. Starting from question Q2, on the properties of disentangled representations that are preserved after transferring, we may conclude Explicitness is usually well maintained, while Modularity and Compactness are reduced as we move from synthetic to real. More precisely, we appreciate a degradation in the global metrics (such as OS and ME), while on the compactness through the analysis of the 1-dimensional pruned representations, we notice that some FoV may transfer very well.
@@
Q1, a much wider question discussing under what circumstances transfer is effective leads us to conclude that some structural similarity between Source and Target datasets is necessary, including similar ranges/granularity of variations of related factors. While real Source data, as long as they are DRL-compliant, might be quite effective in transferring, the results of our study suggest one could design synthetic data to capture specific factors of the problem of interest.
Future directions will target more specific applications, such as biomedical image classification or action recognition from videos, to discuss and relate the general results we are reporting in this paper to more specific and challenging domains.
@@ C.1 Datasets
Fig. 7 reports the main information about the used dataset, e.g. FoV and number of classes, together with some examples of the original dataset, plus some samples of one variant of the given dataset, such as Color-dSprites and Noisy-Color-dSprites. The only exception is Shapes3D which does not have any variant, so different samples drawn from the same dataset are shown.
@@ C.5 GBT & MLP performance distribution
C.6 MLPs performances
@@
C.7 GBT & MLP performances standard deviation

### v3

@@ 5 Conclusions
In this paper, we discussed the potential of transferring a Disentangled Representation as a strategy to address disentanglement in real data. We learned the representation from a Source Dataset in a weakly supervised manner and transfered it to a Target Dataset, where supervision on the FoVs was difficult or impossible to obtain. We identified three main scientific questions, summarised in Section 3.2, which we recall to draw conclusions on our study. Starting from question Q2, on the properties of disentangled representations that are preserved after transferring, we may conclude Explicitness is usually well maintained, while Modularity and Compactness are reduced as we move from synthetic to real. More precisely, we appreciate a degradation in the global metrics (such as OS and ME), while on the compactness through the analysis of the 1-dimensional pruned representations, we notice that some FoV may transfer very well.
@@
Q1, a much wider question discussing under what circumstances transfer is effective, leads us to conclude that some structural similarity between Source and Target datasets is necessary, including similar ranges/granularity of variations of related factors. A quantification of the similarity among datasets is still under investigation; the results of our study suggest one could design synthetic data to capture/disentangle specific factors of interest.
Future directions. Currently, we are exploring quantitative methods to assess the distance between Source and Target datasets. In the near future will target more specific applications, such as biomedical image classification or action recognition from videos, to discuss and relate the general results we are reporting in this paper to more specific and challenging domains.
Acknowledgement
We acknowledge the financial support from PNRR MUR Project PE0000013 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU, CUP J33C24000430007
@@ C.1 Datasets
Table 7 reports the main information about the used dataset, e.g. FoV and number of classes, together with some examples of the original dataset, plus some samples of one variant of the given dataset, such as Color-dSprites and Noisy-Color-dSprites. The only exception is Shapes3D which does not have any variant, so different samples drawn from the same dataset are shown.
@@
[…]
Isaac3D
FoV $\bm{\#}$ values
Object Shape 3
Object Scale 4
Object Color 4
Camera Height 4
Robot X-movement 8
Robot Y-movement 5
Lighting Intensity 4
Lighting Y-direction 6
Wall Color 4
@@ C.5 GBT & MLP performance distribution
C.6 Transfer from Shapes3D to Isaac3D
In this section, we report the results of the transfer from Shapes3D to Isaac3D, see Table 9 for the Explicitness and Table 10 for Modularity and Compactness.
We observe that the transfer is overall effective according to all the disentanglement scores even though the Target dataset is much more complex than the Source. If we compare the results of the transfer from Shapes3D to other real datasets such as Coil100 (Table 4) and RGBD-Objects (Table 5), we notice that in the case of Isaac3D the boost of finetuning is more noticeable.
This suggests that, if source and target have common FoVs with similar appearance we can obtain reasonable performances on a real Target dataset even if the source synthetic dataset is much simpler.
Table 9: Transfer from Shapes3D (Source) to Isaac3D (Target). Average classification accuracy over the 20 models of the GBT classifier, before and after fine-tuning (see Table 2).
Mean accuracy on FoVs(%)
Pruned Object shape (3) Object scale (4) Camera height (4) X-movement (8) Y-movement (5) Light intensity (4) Light y-direction (6) Object color (4) Wall color (4) All
34.9 (+5.1) 54.0 (+34.6) 39.2 (+17.4) 33.9 (+29.6) 23.0 (+4.3) 83.6 (+14.0) 85.4 (+12.7) 29.8 (+13.9) 78.1 (+18.9) 51.3 (+16.7)
33.8 (+3.6) 40.1 (+24.1) 33.5 (+11.1) 24.7 (+15.3) 21.6 (+2.7) 69.4 (+17.3) 67.5 (+14.4) 27.0 (+7.3) 61.5 (+16.6) 42.1 (+12.5)
Table 10: Transfer from Shapes3D (Source) to Isaac3D (Target). The Compactness and Modularity scores of the same models of Table 9.
Modularity(%) Compactness(%)
Pruned Our (OS) DCI Our (MES) MIG
25.1 (+9.7) 6.3 (+16.1) 21.2 (+10.2) 2.2 (+5.9)
C.7 MLPs performances
@@
C.8 GBT & MLP performances standard deviation